OpenStudy (anonymous):

find integral of \frac{4}{3} \int\limits_{3}^{5}\sqrt{x^2 - 9}

OpenStudy (anonymous):

\[\frac{4}{3} \int\limits\limits_{3}^{5}\sqrt{x^2 - 9} \]

OpenStudy (agent0smith):

Have you done trig substitutions? http://www.integralcalc.com/wp-content/uploads/2012/10/sin-tan-sec.png

OpenStudy (anonymous):

it is hyperbolic sub. but im done well knowledgeable about how to execute the integral form there

OpenStudy (anonymous):

not well knowledgeable* from*

OpenStudy (agent0smith):

Ah okay. I haven't done those in a while, but http://mathworld.wolfram.com/HyperbolicSubstitution.html and http://tutorial.math.lamar.edu/Classes/CalcI/DiffHyperTrigFcns_files/eq0007MP.gif so let x = 3coshu

OpenStudy (agent0smith):

\[\large x = 3\cosh u\] \[\large 1 = 3\sinh u \frac{du}{dx}\] \[\large dx = 3 \sinh u du\] \[\large \int\limits \sqrt{x^2 - 9} dx = \int\limits\limits\limits \sqrt{9 \cosh^2 u - 9} ( 3 \sinh u) du\] (ignoring all the fractions and limits)

OpenStudy (agent0smith):

\[\large \int\limits\limits\limits\limits \sqrt{9( \cosh^2 u - 1)} ( 3 \sinh u) du \] \[\large 9\int\limits \sqrt{\sinh^2 u} ( \sinh u) du = 9 \int\limits \sinh^2 u du \] Now you can prob integrate by changing the sinhu back to it's exponential form http://tutorial.math.lamar.edu/Classes/CalcI/DiffHyperTrigFcns_files/eq0003MP.gif